OpenStudy (anonymous):

Jared has four pieces of rope that are each 12.75 feet long. What is the total length of rope he has? 51 feet 50.8 feet 12.79 feet 16.75 feet

OpenStudy (anonymous):

add all the length = 12.75 +12.75+12.75+12.75 = 4*12.75

OpenStudy (anonymous):

51

OpenStudy (anonymous):

All you have to do is times 12.75 and 4......